OpenStudy (anonymous):

Will Give Medal!!! Amy and Elizabeth both work in sales. Each week Amy earns $18/h plus a $40 bonus, and Elizabeth earns $12/h plus a $100 bonus How many hours must Amy and Elizabeth each work in a week in order to earn the same amount of money? For that many hours, how much money would each earn? They would each have to work ___hours, and they would earn $___

OpenStudy (anonymous):

Well what you can do is right each of their earnings as equations with an unknown variable in the form of the number of hours. So we know that Amy earns $18 per hour and Elizabeth $12 per hour and that each has a bonus that we can add on as a constant. So the equations for each of their earnings in a week would be: Amy: 18h + 40 Elizabeth: 12h +100. You want these two to be the same so put them equal to each other. 18h + 40 = 12h + 100 Rearranging gives: 6h = 60 and then you can find h.

OpenStudy (anonymous):

To find the amount that they have earned in that time, just put your value for h back into either of the original equations. So 18x10 + 40 = 220 or 12x10 + 100 = 220.

OpenStudy (anonymous):

so, They work 10 hours..? and somehow I got 1 for the number of money they get.. im sorry im really bad at math .-. @alrightatmaths

OpenStudy (anonymous):

Exactly, they work 10 hours and get the same amount of money. To find the amount of money, just put h=10 into one of the first two equations. If you take 18h +40, just put 10 in the place of h. 18x10 +40. 18x10 = 180 and 180+40 = 220.

OpenStudy (anonymous):

SO in conclusion.. lol its either really 1 or 220 >.< haha

OpenStudy (anonymous):

Lol I understand now , Thanks so much! @alrightatmaths